OpenStudy (rational):

I don't get this proof : prove that number of divisors of any positive number \(\large n\) is \(\large \le 2\sqrt{n}\)

OpenStudy (rational):

edited the question, please see again :)

OpenStudy (rational):

Is there any way to derive this usign divisors formula : number of divisors of \(n = p_1^{k_1}p_2^{k_2}\cdots p_r^{k_r} \) is \((k_1+1)(k_2+1)\cdots (k_r+1)\)

OpenStudy (tkhunny):

Divisors come like this... a*b = The number we seek, n. Without loss of generality, let's say \(a \le b\). This leaves a = b for n a perfect square. Let's just think about the number 'a'. What are it's possibilities. 1) It is limited to \(a \le \sqrt{n}\) 2) It is limited by \(Count(a) \ge Prime\;Numbers\;less\;than\;(or\;equal\;to) \sqrt{n}\) Your unique prime factorization is a great way to go. How may ways are there to shuffle those things?

OpenStudy (rational):

\(\large a \le b\) gives half (more or less) of the divisors, right ?

OpenStudy (anonymous):

@tkhunny I haven't seen the count function outside of Excel... Does it behave at all like the floor function?

OpenStudy (rational):

\[\large n = d*\dfrac{n}{d}\]

OpenStudy (ikram002p):

do you have any proof ? cuz im thinking of something like this \(\sqrt n \le n! \) its only a though >.< i wanna see a proof

OpenStudy (ikram002p):

so if we assumed n is composite , then we cheking the factors \(\le \sqrt n\) The Sieve of Erathosthens >.<

OpenStudy (rational):

how is it related to sieve of eratosthenes ? we're checking for divisors, not primes right ?

OpenStudy (ikram002p):

well first we know that primes factors \(\le \sqrt n\) right ? and any other factor is a formed from the factors hmm but i still need how to onclude the 2 part

OpenStudy (rational):

It seems the upper limit is not strong, is there any number which has divisors = \( 2\sqrt{n}\) ?

OpenStudy (ikram002p):

im not convinced with the upper thingy , my point is how many combination of the primes factors we could get less than n

OpenStudy (rational):

Oh i get what you're trying

OpenStudy (rational):

did you understand the proof in MSE ?

OpenStudy (ikram002p):

if we have \(p_1,p_2,p_3,.., p_i\) ( primes factors ) then the number of primes factors is d\(\le \sqrt n\) now :- \(p_i,2p_i,3p_i,.., kp_i \) \(p_i,p_i^2,p_i^3,.., p_i^c \) such that \(kp_i <n , p_i^c <n\) are all factors hmmm

OpenStudy (tkhunny):

The exact number of divisors is expressed as: If \(s = k_{1} + k_{2} + k_{3} + ... + k_{r}\) Then \({s}\choose{0}\)+\({s}\choose{1}\)+\({s}\choose{2}\)+...+\({s}\choose{s}\)

OpenStudy (rational):

exactly ! number of divisors is a combination - number of different ways of choosing exponent of a prime number in the prime factorization

OpenStudy (ikram002p):

rash , i did understand but i dont understand why http://prntscr.com/4f4k51

OpenStudy (ikram002p):

yeah :o

OpenStudy (rational):

ikram : seive of eratosthenes gives you number of prime factors are \(\le \sqrt{n}\) I don't get your statements after this, could you please elaborate ? :)

OpenStudy (ikram002p):

this is what i meant to do xD number of divisors is a combination - number of different ways of choosing exponent of a prime number in the prime factorization but im not neat to say that >.<

OpenStudy (rational):

http://prntscr.com/4f4k51 i don't understand that either, they're claiming stuff without explaining

OpenStudy (ikram002p):

ok so the question would be like this if u have a set of primes ( that are also factors of n )\(A\) of size \(\sqrt n\) , how many factors of n you could form from A . (or something like this )

OpenStudy (ikram002p):

i think we need the floor function series here :o

OpenStudy (ikram002p):

:3 not good with permentation let A=(p_1,p_2,p_3 ,....,p_k) factors set =\( ( p_1,p_2,p_3 ,....,p_k , p_1^2,p_2^2,p_3^2\\ ,....,p_k^2 ,....p_1^r,p_2^r,p_3^r ,....,p_k^r , p_1p_2,p_1p_3 ,....,p_1p_k.... ect )\)

OpenStudy (ikram002p):

but with condition that any element <n :3

OpenStudy (rational):

see if below makes sense : \[\large n = d*\dfrac{n}{d}\] Clearly either \(d\) or \(\dfrac{n}{d}\) is \(\large \le n\), 1) \(\color{Red}{1\le d \le \sqrt{n}} \implies n \ge \dfrac{n}{d} \ge \sqrt{n}\) 2) \(\sqrt{n} \le d \le n \implies \color{red}{ \sqrt{n} \ge \dfrac{n}{d} \ge 1}\)

OpenStudy (rational):

that gives : 1) there can be atmost \(\color{Red}{\sqrt{n}}\) values possible for \(d\) less than or equal to \(\color{Red}{\sqrt{n}}\) 2) there can be atmost \(\color{Red}{\sqrt{n}}\) values possible for\( \dfrac{n}{d}\) less than or equal to \(\color{Red}{\sqrt{n}}\)

OpenStudy (ikram002p):

it make sense now :o

OpenStudy (rational):

So, the upper limit for total number of divisors would be : \(\large \color{red}{\sqrt{n} + \sqrt{n}}\)

OpenStudy (ikram002p):

="(

OpenStudy (ikram002p):

that was ... >.<

OpenStudy (rational):

same here, makes some sense now :) did u get how \(\large \tau (n) = \prod \limits_{1\le i \le r} (k_i+1)\) ?

OpenStudy (ikram002p):

this one i know why

OpenStudy (ikram002p):

but could u proof it , using this ?

OpenStudy (rational):

http://prntscr.com/4f4xbf

OpenStudy (ikram002p):

no , i know this >,< ok nvm that was not my question

OpenStudy (rational):

what was your question?

OpenStudy (ikram002p):

do you have another proof for d<= 2 sqrt n ?

OpenStudy (tkhunny):

So also, \(1 + 2^{s}\). Quite a bit simpler than the previous expression. Of course, expressing this in terms of 'n' is the real matter at hand.

OpenStudy (rational):

I think it equals \(\large 2^s\) when \(\large n \) is square free : \(\large p_1^1 p_2^1 \cdots p_s^1\)

OpenStudy (ikram002p):

?

OpenStudy (rational):

thanks everyone ! you would never know how much you have helped me !!

OpenStudy (ikram002p):

u solve it by ur own -_-

OpenStudy (rational):

if \(\large n\) is square free, \(\large n\) has no prime exponent greater than 1, so in the prime factorization, each prime has \(\large 2\) choices

OpenStudy (rational):

if \(n\) is a product of \(s\) distinct primes, then total number of divisors would be : \(\large 2\times 2\times \cdots \text{s times} = 2^s\)

OpenStudy (ikram002p):

yeah

OpenStudy (rational):

lets work one more last question for the day new thread

OpenStudy (ikram002p):

hope its simple >.<